Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 94     Entry time: 22 Mar 2023, 21:49
Author: Adam Round 
Group: SPB 
Subject: Shift end 

X-ray delivery

  • 9 keV 1 to 165 pulses ~ 300 mj per pulse - stable delivery during the day

Optical laser delivery

  • ns lasers for illumination only

Achievements/Observations

  • Focus alignment
  • found illumination / shield (100 micron) conditions to optimise background on detector
  • observed that the flourecence signal on the diode was very low and this appears corelated with drift in the focus - after focus optimisation signal was again visible
  • several runs collected with varying attenuation with the aim to try to corelate flourecence signal on diode with beam intensity

Issues

  • Too high background from sample at full transmission required exchange of copper shield form 50 to 100 microns
  • disc detaching from motor spindle (twice) - first time disc blocked gate valve, second bolt and washer are in the gate valve mechanisme space
  • in both cases detector had to be warmed to allow venting.
ELOG V3.1.4-7c3fd00